Circular No. 7657

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001cz IN NGC 4679 R. Chassagne, Ste. Clotilde, Île de Réunion, reports the discovery of

an apparent supernova (mag ∼ 14.7) on unfiltered CCD images taken on July 4.64 and 5.61 UT with a 0.30-m reflector. SN 2001cz is located at α = 12^{h}47^{m}30^{s}.17, δ = −39^{o}34′48′′.1 (equinox 2000.0), which is 0′′.6 west and 31′′.4 south of the nucleus of NGC 4679. Images taken on 2000 Mar. 10.9, and on 2001 Mar. 21.9 and May 21.7, show nothing at this position (limiting mag ∼ 19.0), and SN 2001cz does not appear on a Digital Sky Survey (U.K. Schmidt) image taken on 1982 May 16.47 (limiting mag ∼ 21.0).

SUPERNOVAE 1993J AND 1979C P. Chandra, Tata Institute of Fundamental Research (TIFR), Mumbai;

S. Bhatnagar, National Centre for Radio Astrophysics, TIFR, Pune; and A. Ray, TIFR, Mumbai, report the detection of SN 1993J in NGC 3031 for the first time at 610 MHz with the Giant Meter Radio Telescope (GMRT) on 2001 Mar. 24. The flux density at 609.37 MHz was 40.9 ± 1.2 mJy. The host galaxy’s center had a flux density of 70.3 mJy. SN 1993J was also detected at 1420 MHz with the GMRT on 2001 June 2 (32.2 ± 0.3 mJy) and on 2000 Dec. 16 (47.7 ± 0.6 mJy). They also report the first detection of SN 1979C in NGC 4321 with the GMRT at 610 MHz on 2001 Mar. 24 with a flux density of 13.1 ± 1.4 mJy. The host galaxy’s center had a flux density of 58.4 mJy. SN 1979C was also detected at 1420 MHz band by the GMRT on 2000 Nov. 9 (36.6 ± 0.6 mJy).

2001 KX_{76}

R. L. Millis, Lowell Observatory, on behalf of the Deep Ecliptic Survey Team, reports the following photometric magnitudes of 2001 KX_{76} (cf. MPEC 2001-N01) obtained by J. L. Elliot, S. D. Kern, and D. J. Osip at the 6.5-m Magellan Telescope on June 9: R = 19.6 ± 0.1, B −V = +0.95 ± 0.15, V −R = +0.5 ± 0.1, V −I = +1.0 ± 0.1. Best-fitting orbital elements by L. Wasserman place 2001 KX_{76} at r = 43.5 ± 0.2 AU and Δ = 42.5 ± 0.2 AU; if these distances are correct, 2001 KX_{76} appears to be intrinsically brighter than (20000) Varuna = 2000 WR_{106} (IAUC 7554). Radiometric, astrometric, and spectroscopic observations of 2001 KX_{76} are encouraged.

2001 July 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green